A ray of light along x+(3y)^1/2 =3 gets reflected upon reaching x-axis, the equation of the reflected rays is

Grade:upto college level

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Equation of ray of light:
x + \sqrt{3}y = 3
\sqrt{3}y = -x + 3
y = -\frac{1}{\sqrt{3}}x + \sqrt{3}
This line make 150 degree with +ve x – axis.
So this makes 60 degree with normal.
So reflected light make 30 degree with +ve x – axis.
Equation of reflected light:
(y - \sqrt{3}) = \frac{1}{\sqrt{3}}(x - 0)
y = \frac{x}{\sqrt{3}} + 3